# Hierarchical Hybrid Learning for Long-Horizon Contact-Rich Robotic Assembly

### v1

@@ IV-B High-level Policy: Composing Primitives via Imitation
Figure 4: Pose Estimation Examples. Our method gives accurate pose estimation of differently shaped parts. The top row shows the input images and the bottom row shows the pose estimation. Red, green, blue colors indicate the $xyz$ axes in the canonical space, respectively.
IV-B1 Data Collection
As our low-level primitives are executed with either RL policies or MP policies, no data collection is needed. In this section, we focus on the data collection process [^1] Although datasets like FMB [28] exist, our setup and hardware differ. We use a UR10e instead of the Panda robot. Additionally, we found that images captured by the RealSense camera were not accurate enough for pose estimation, so we opted for the Zivid camera. As a result, we need to collect our own data to suit our setup. for the high-level policy. After training the individual action primitives, a human expert sequentially selects them to complete the multi-stage peg insertion task, using a keyboard. The continuous input parameter of each action primitive is the pose obtained from pose estimation. For example, to initiate the grasp primitive, the demonstrator selects the primitive index, and the object pose obtained from the pose estimation module is passed to the motion planner to execute the primitive. To enhance robustness, we augment the dataset by introducing noise into the robot’s state. The dataset denoted as $\mathcal{D}=\{(\mathbf{s},\mathbf{a},\mathbf{u})\}$, consists of sensor observations $\mathbf{s}$, primitive indexes $\mathbf{a}$, and corresponding continuous primitive parameters $\mathbf{x}$.
@@ IV-B1 Data Collection
The dataset consists of two types of demonstrations: “successful” and “recovery” trials. In approximately 20 trials, the demonstrator successfully inserts the object, while in the remaining 20 trials, failures occur, followed by recovery actions demonstrated by the operator. These recovery trials are crucial to making the learned policy more robust to errors.
@@
In summary, we allow the demonstrator to select the discrete primitive to execute while the continuous input parameter comes from pose estimation. This is a novel way to collect demonstration data as it is often difficult and time-consuming to collect teleoperated demonstrations for high-precision tasks.
IV-B2 Pose Estimation
@@
As shown in Fig. 2, object pose is needed by the high-level policy and also serves as the input parameter to the low-level primitive skills [43, 49, 15], such as grasp. We integrate the state-of-the-art pose estimation method CPPF++ [50, 51], which demonstrates strong generalization from simulation to real-world scenarios. Specifically, given a 2D RGB-D image $\mathcal{I}$, the model produces an accurate 6D pose $\xi\in\mathfrak{se}(3)$, where the first three components represent rotation and the last three correspond to translation. While the original CPPF++ is designed for category-level pose estimation, we adapt it for instance-level pose estimation by using distinct industrial CAD models as different “categories”. To enhance precision, inspired by the Iterative Closest Point (ICP) [3] algorithm, we propose a post-optimization procedure conducted in the tangent space of the Lie group.
@@
where $p^{*}_{j}$ is the $j$-th point on the back-projected point cloud obtained from the predicted masks, $p_{i}$ is the $i$-th point on the object mesh, and $T_{\xi}$ is the rigid transformation that aligns the canonical object point with the scene, defined as $T_{\xi}(p_{i})=R\cdot p_{i}+t$. We utilize the one-way Chamfer distance due to self-occlusion in the observation and optimize $\xi$ iteratively using the Lietorch [46] library to obtain a refined pose.
Fig. 4 presents some qualitative pose detection results. The high accuracy of the pose estimator enables us to have a high success rate in terms of both the low-level MP policies and the high-level policy.
V Experiments
V-A Task Description
@@
We focus on the multi-stage assembly tasks [28]. As shown in Fig. 1, 9 objects of different shapes must be inserted into the board. The robot must grasp the object, may need to reorient and regrasp it using the fixture depending on its grasp pose, and then insert it into the board. We assume CAD models of all the objects are available and the location of the fixture and the board is known. The objects range from simple or symmetrical shapes, such as rectangle, circle, and oval, to more complex ones, including star, 3-prong, square-circle, and arch. We evaluate methods across subgoal and long-horizon task categories and examine generalization using unseen objects.
V-B Setup and Implementation
The workcell setup includes a UR10e robotic arm paired with a Robotiq 2F-85 gripper and a third-person view Zivid 2+ M60 camera (Fig. 1). The simulation platform for RL training is built on the IsaacLab engine [34], while motion-level manipulation plans are computed using the lazyPRM approach [5]. Model training is performed on a machine equipped with an Intel 3.80GHz i7-10700k CPU and a GeForce RTX 4090 Ti GPU, running on a Ubuntu system.
For pose estimation, we use $512\times 512$ RGB-D image as input. For high-level policy $\pi_{\theta}$, We adopt the Diffusion Transformer model [37] for predicting action feasibility score. Our RL primitive is trained using 1,000 parallel environments in IsaacLab. The maximum task horizon is 25,000 action steps, equivalent to 200 seconds of robot execution at a 125Hz control frequency.
V-C Baselines
End-to-end Learning
@@ V-C Baselines
We evaluate two end-to-end methods: 1) End-to-end IL: Diffusion Policy (DP) [8], a goal-conditioned imitation learning framework that leverages a diffusion model for generating diverse action trajectories. Teleoperated demonstration data is collected in the real world. 2) End-to-end RL: PPO [38], a widely-used RL method that optimizes policy performance through proximal updates, balancing exploration and exploitation while ensuring stable training. Training is conducted in the IsaacLab simulation environment.
Hierarchical Methods
@@
We evaluate two hierarchical methods: 1) MimicPlay [47] learns high-level latent plans from human play data to guide low-level visuomotor control. Teleoperated demonstration data is collected in the real world. 2) Luo et al. [27] propose a Multi-Stage Cable Routing approach through hierarchical imitation learning. Teleoperated demonstration data for both low-level and high-level policies is collected in the real world. We adapt the above methods to our task.
@@
TABLE I: Success rate of the multi-stage task of Hexagon assembly, comparing our method with baseline methods after training on $40$ demonstrations on the Hexagon object.
Method SR (%) $\uparrow$ SPL $\uparrow$
E2E RL [38] 0 0.00 $\pm$ 0.00
E2E IL [8] 0 0.00 $\pm$ 0.00
MimicPlay [47] 40 0.78 $\pm$ 0.05
Luo et al. [27] 50 0.83 $\pm$ 0.13
ARCH (Ours) 55 0.93 $\pm$ 0.11
Human Oracle [28] 65 0.95 $\pm$ 0.08
@@
TABLE II: Success rate for multi-stage assembly task and single-stage tasks by object. Our system demonstrates the ability to generalize to unseen objects.
@@
Figure 5: Data efficiency comparison across imitation learning methods. Trained with just 10 demonstrations, our system achieves a higher success rate compared to baselines, which require significantly more data to perform well.
V-D Evaluation Metrics
For each trial, the selected object is placed within the grasp randomization region, centered at one of four random positions, and oriented in one of five possible angles. A total of 20 trials are conducted. For the grasping primitive, the object must be securely gripped by the gripper to prevent it from falling after being lifted. Finally, the object’s bottom surface must be fully inserted into the corresponding hole.
The primary metric is the task success rate (%). We found that failures primarily stem from two challenging primitives: grasping and insertion. To better understand the sources of error, we introduce two specific metrics: ”Grasped” and ”Inserted,” which measure the success rates of these two primitives.
Success Rate (SR)
the percentage of successfully completed multi-stage tasks (instead of each primitive) out of the total multi-stage tasks attempted.
SPL
Success weighted by Path Length, as Eq. 5.
$\textrm{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\cdot\frac{l_{i}}{\max(l_{i},l^{opt}_{i})},$ (5)
where $N$ is the total number of evaluation trials, $S_{i}$ is 1 if the $i$-th trial is successful and 0 otherwise, $l^{opt}_{i}$ is the shortest path (number of primitives) to the goal, and $l_{i}$ is the actual path length taken by the agent.
% Grasped
the percentage of target objects that are successfully grasped as a single-stage task.
% Inserted
the percentage of target objects that are successfully inserted as a single-stage task.
V-E Experiment Analysis
@@
As shown in Tab. I, long-horizon tasks often involve sparse rewards, which complicates RL training. Consequently, end-to-end (E2E) RL [38] fail at such long-horizon assembly tasks. With diffusion models, E2E IL [8] falls short in handling high-precision tasks and requires a large number of expert demonstrations. The hierarchical baselines, MimicPlay [47] and Luo et al. [27], achieve better success rate than the E2E methods, but struggle with grasp failures and consequently have lower overall success rate than our method. Our ARCH demonstrates higher success rates and SPL compared to the baselines due to our hierarchical hybrid design. By combining model-based and adaptive learning-based components, our approach achieves both high precision and flexibility. Although our RL insertion primitive is trained solely in simulation on the Hexagon object for the insertion primitive, it transfers well to real-world scenarios and to other objects. Finally, Human Oracle is not a baseline but serves as the upper bound for high-level policy. The human oracle selects primitives with near-flawless accuracy and can self-correct mistakes. Failures with this method are primarily due to pose estimation or grasp errors.
Generalization to Unseen Objects
@@
Our system has been trained exclusively with the Hexagon object, both for the RL primitive and for the high-level policy. Tab. II shows that our system generalizes well to unseen objects without any fine-tuning. Objects with a narrow edge, such as the SquareCircle and DoubleSquare, present more challenges during the grasping stage. Additionally, certain object shapes, such as the Star, SquareCircle, and Arch, present more challenges to pose estimation.
Data Efficiency
As Fig. 5 shows, ARCH demonstrates high data efficiency. Collecting teleoperated demonstration data is both costly and time-consuming, making it crucial to maximize the utility of each sample. Our approach outperforms previous methods in terms of sample efficiency, enabling the system to achieve effective learning with fewer training examples. Additionally, our system excels in multi-stage task success rates, indicating that it can maintain performance over extended sequences of actions. This capability not only reduces the overall data collection burden but also enhances the practical applicability of our method in real-world scenarios, where data acquisition can be a limiting factor.
Robustness
@@
Our high-level policy exhibits robustness to human disturbances and facilitates failure recovery, allowing the robot to recover from unexpected situations absent in the demonstration data. For example, we observed that when the grasp primitive fails, the policy automatically triggers another grasp attempt using the new pose estimation.
Failure Cases
Although our hierarchical system performs well, long-horizon assembly remains challenging due to several factors. We have identified that failures are primarily caused by errors in grasp pose estimation and RL insertion primitive.
The sources of error include:
  • Sensor: Particularly the depth camera, which impacts object/grasp pose estimation.
  • Grasp Pose Estimation: Errors in estimating the precise position and orientation of the object and grasp can lead to failure initially and impact the overall assembly process.
  • RL: Inherent uncertainties in the RL policy.
VI Conclusion, Limitations, and Future Work
@@
Although our system demonstrates excellent performance and generalizability, there are limitations. For example, we manually identify the object beforehand. Incorporating an object detection method would be beneficial in the future. We would also like to extend to other types of assembly tasks and incorporate a more comprehensive primitive skill library.

### v2

@@ 3.2 High-level Policy: Composing Primitives via Imitation
3.3 Data Collection
As our low-level primitives are implemented with either RL policies or MP policies, no demonstration data is needed. In this section, we focus on the data collection process for the high-level IL policy. Our hierarchical framework abstracts away the motion details from the demonstrator and allows them to demonstrate by selecting primitives and parameters instead of teleoperation. More specifically, from the library of pre-defined action primitives (Section 3.1), a human demonstrator uses a keyboard to select the most appropriate primitive and its parameters to complete the assembly task at hand. For example, to initiate the GRASP primitive of object $o$, the demonstrator selects the primitive index and the object index. The pose estimation module (Section 3.4) estimates the object pose $p_{o}$ and the continuous parameter, pre-grasp pose $p_{g}$, is calculated by adding a fixed offset from $p_{o}$. $p_{g}$ is then passed to the motion planner to execute the primitive. To initiate the INSERT primitive of object $o$, the demonstrator selects the primitive index and the object index. As described in the assumptions, the insertion goal pose $g$ for each object is known with a small margin of error. A pre-trained RL policy is used with $g$ as its goal pose.
@@ 3.3 Data Collection
The dataset denoted as $\mathcal{D}=\{(\mathbf{s},\mathbf{a},\mathbf{x})\}$, consists of sensor observations $\mathbf{s}$, primitive indexes $\mathbf{a}$, and corresponding primitive parameters $\mathbf{x}$. The dataset consists of two types of demonstrations: “successful” and “recovery” trials. In half of the trials, the demonstrator successfully inserts the object, while in the other half, failures occur, followed by recovery actions demonstrated by the operator. These recovery trials are crucial to making the learned policy more robust to errors.
@@
In summary, we allow the demonstrator to select the discrete primitive and parameter to execute while the continuous input parameter comes from pose estimation. This is a novel way to collect demonstration data as it is often difficult and time-consuming to collect teleoperated demonstrations for high-precision tasks.
3.4 Pose Estimation
@@
As shown in Figure 2, object pose is needed by the high-level policy and also serves as the input parameter to the low-level primitive skills, such as grasp. We integrate the pose estimation method CPPF++ [55, 56], which demonstrates strong generalization from simulation to real-world scenarios. Specifically, given a 2D RGB-D image $\mathcal{I}$, the model produces an accurate 6D pose $\xi\in\mathfrak{se}(3)$, where the first three components represent rotation and the last three correspond to translation. While the original CPPF++ is designed for category-level pose estimation, we adapt it for instance-level pose estimation by using distinct CAD models as different “categories”. To enhance precision, inspired by the Iterative Closest Point (ICP) [57] algorithm, we propose a post-optimization procedure conducted in the tangent space of the Lie group.
@@
where $p^{*}_{j}$ is the $j$-th point on the back-projected point cloud obtained from the predicted masks, $p_{i}$ is the $i$-th point on the object mesh, and $T_{\xi}$ is the rigid transformation that aligns the canonical object point with the scene, defined as $T_{\xi}(p_{i})=R\cdot p_{i}+t$. We utilize the one-way Chamfer distance due to self-occlusion in the observation and optimize $\xi$ iteratively using the Lietorch [58] library to obtain a refined pose. Figure 4 in the Appendix presents some qualitative pose detection results. The high accuracy of the pose estimator enables us to have a high success rate in terms of both the low-level MP policies and the high-level policy.
4 Experiments
4.1 Task Description and Setup
Recent advancements in robotic learning have also highlighted the need for benchmarks that effectively balance generalization and the complexity of manipulation tasks. The Functional Manipulation Benchmark (FMB) [59] aims to bridge this gap by defining functional manipulation as a series of relevant behaviors, such as grasping, repositioning, and physically interacting with objects. Building on FMB, we want to extend its emphasis on contact-rich dynamics and object diversity. For these reasons, we use the multi-peg assembly task introduced in FMB, along with two more challenging custom-designed multi-part assembly tasks, to showcase the generalizability of our approach.
@@ 4.1 Task Description and Setup
We focus on three assembly tasks: FMB Multi-peg Assembly in [59], 5-part Beam Assembly, and 9-part Stool Assembly, as shown in Figure 1. FMB Assembly is tested on the real robot, whereas the other two assemblies are in simulation only. For FMB Assembly, 9 objects of different shapes must be inserted into the board. The robot must grasp the object that is randomly placed on the table, may need to reorient and regrasp it using the fixture depending on its initial pose, and then insert it into the board. More task and setup details can be found in the Appendix. We evaluate single-step and long-horizon categories and examine generalization to unseen objects and tasks.
4.2 Baselines
@@
We evaluate two end-to-end (E2E) baselines: 1) E2E IL: Diffusion Policy (DP) [9], a goal-conditioned imitation learning framework that leverages a diffusion model for generating diverse action trajectories. Teleoperated demonstration data is collected in the real world. 2) E2E RL: PPO [52], a widely-used RL method that optimizes policy performance through proximal updates, balancing exploration and exploitation while ensuring stable training. Training is conducted in the IsaacLab simulation environment [51].
@@
We also evaluate two hierarchical baselines: 1) MimicPlay [5] learns high-level latent plans from human play data to guide low-level visuomotor control. Teleoperated demonstration data is collected in the real world. 2) Luo et al. [49] propose a Multi-Stage Cable Routing approach through hierarchical imitation learning. Teleoperated demonstration data for both low-level and high-level policies is collected in the real world. We adapt the above methods to our task.
@@
Table 1: Success rate SR(%) of the multi-stage FMB Assembly of Hexagon, Beam Assembly, and Stool Assembly, comparing our method with baseline methods after training with 1010 demonstrations, with the human oracle being the upper bound.
Methods FMB (real) Beam (sim) Stool (sim)
Metrics SR (%) $\uparrow$ SPL $\uparrow$ SR (%) $\uparrow$ SPL $\uparrow$ SR (%) $\uparrow$ SPL $\uparrow$
E2E RL [52] 0 0.00 0 0.00 0 0.00
E2E IL [9] 0 0.00 0 0.00 0 0.00
MimicPlay [5] 20 0.16 15 0.12 10 0.08
Luo et al. [49] 25 0.19 15 0.14 20 0.18
ARCH (Ours) 55 0.51 55 0.51 45 0.44
Human Oracle [59] 65 0.55 60 0.62 50 0.49
4.3 Evaluation Metrics
A total of 20 trials are conducted for each task. For each trial, the selected object is initialized randomly in the grasp region. For the GRASP primitive to succeed, the object must be securely gripped by the gripper without falling after being lifted. For the INSERT primitive to succeed, the object must be fully inserted into the corresponding receptacle.
The main metric is the overall task success rate (%). We found that failure primarily stems from grasping and insertion. Hence, we also introduce two metrics: ”% Grasped” and ”% Inserted”, to measure the success rates of these two primitives.
Success Rate (SR%): percentage of successful completion of the long-horizon tasks.
Success weighted by Path Length (SPL):
$\textrm{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\cdot\frac{l_{i}}{\max(p_{i},l_{i})},$ (4)
where $N$ is the total number of trials, $S_{i}$ is a binary indicator of success in trial $i$, $l_{i}$ is the shortest path (number of primitives) to the goal, and $p_{i}$ is the actual path length taken by the agent.
% Grasped: percentage of successful grasps as a single-stage task.
% Inserted: percentage of successful insertions as a single-stage task.
@@ 4.3 Evaluation Metrics
Table 2: Success rate SR(%) of the multi-stage FMB Assembly task after training with 1010 demonstrations and success rate of the single stages by object. Our system demonstrates the ability to generalize to unseen objects.
@@
4.4 Experiment Analysis
@@
Comparative Performance Analysis. First, we compare with end-to-end baselines and hierarchical approaches. For methods that utilize a low-level primitive library, we ensure they remain identical across all methods to ensure a fair comparison. As shown in Table 1, long-horizon tasks often involve sparse rewards, which complicates RL training. Consequently, E2E RL [52] fails at such long-horizon assembly tasks. With diffusion models, E2E IL [9] falls short in handling high-precision tasks and requires a large number of expert demonstrations. The hierarchical baselines, MimicPlay [5] and Luo et al. [49], achieve better success rate than the E2E methods, but struggle with contact-rich steps such as insertion due to the small number of demonstrations available. Finally, Human Oracle is not a baseline but serves as the upper bound for high-level policy. The human oracle selects primitives with near-flawless accuracy and can self-correct mistakes. Failures with this method are due to pose estimation and grasp errors. ARCH demonstrates significantly higher success rates than baseline methods due to our hierarchical framework and the hybrid MP-RL design of the low-level library. The advantage of ARCH becomes more pronounced on the more challenging Beam Assembly and Stool Assembly tasks compared to baseline methods. By combining model-based and learning-based components, our approach achieves both high precision and flexibility. Failures with our approach are mainly due to inaccuracies in pose estimation and limitations in the RL policies. Since pose estimation and insertion policies are extensively studied areas, numerous established methods [60, 8] exist to enhance performance; however, exploring these improvements is beyond the scope of this work.
@@
Generalization to Unseen Objects.   Our system has been trained exclusively with the Hexagon object, both for the RL primitive and for the high-level policy. Table 2 shows that our system generalizes well to unseen objects without any fine-tuning. Objects with a narrow edge, such as the SquareCircle and DoubleSquare, present more challenge during the grasping stage. Additionally, certain object shapes, such as the Star, SquareCircle, and Arch, present more challenge to pose estimation.
Data Efficiency.   As Table 1 shows, ARCH achieves high success rates with merely $10$ demonstrations, whereas baseline methods would require significantly more data. This capability not only reduces the overall data collection burden but also enhances the practical applicability of our method in real-world scenarios, where data acquisition can be a limiting factor.
@@
Robustness.   Our high-level policy exhibits robustness to single-stage failures and human disturbances via retries. For example, we observed that when the grasp primitive fails, the policy automatically triggers another grasp attempt using the new pose estimate. This failure recovery capability contributes to the higher success rate of ARCH.
5 Conclusion
@@
6 Limitations
Cross-Task Generalization. Although the learned high-level policy transfers across novel objects within a given task, it remains task-specific: separate training is still required when moving from, for example, FMB Assembly to Stool Assembly. However, training a high-level policy for a new task requires only a handful of demonstrations and no teleoperation, keeping the adaptation effort minimal.
Object Identification Pipeline. All objects are manually labeled and indexed prior to execution. Replacing this offline step with a perception module—e.g., a vision-language model–based detector that auto-assigns semantic IDs—would streamline this process and improve overall efficiency.
Scalability and Generality of the Primitive Library. Our approach assumes that tasks can be decomposed into a finite set of parameterized primitives. This assumption may pose scalability challenges for tasks that require fundamentally new skills. However, in practice, we observe that a relatively small library of primitives—automatically discovered from demonstrations—generalizes effectively across a wide range of long-horizon tasks. Moreover, in industrial settings, tasks generally require only a small set of primitives. In addition, our framework is modular and can readily integrate newly introduced primitives when needed, making it naturally extensible to novel tasks.
Appendix A Appendix
A.1 Task Descriptions
In addition to FMB Assembly, we introduce two more challenging long-horizon assembly tasks, both of which involve multi-stage dependencies. These dependencies make the tasks more difficult—for instance, the feet can only be inserted after the legs are correctly assembled. If the legs are not properly inserted, subsequent steps will fail.
FMB Assembly: This task involves assembling 9 distinct objects with varying shapes. For each object, the robot must first grasp it, then perform a sequence of reorientation actions using an environment fixture, followed by an insertion into the board. The time horizon for each object ranges from 20 to 40 seconds.
5-part Beam Assembly: The base part is fixed to the table. The task involves inserting two legs into the base, followed by inserting two feet onto the legs. The time horizon for each object ranges from 20 to 40 seconds.
9-part Stool Assembly: As shown in Figure 3, the base is also fixed to the table. This task involves inserting four legs into the base, and then inserting four feet onto the legs. The increased number of objects and stages makes this task significantly more challenging and long-horizon. The time horizon for each object ranges from 20 to 40 seconds.
For visualizations of the initial and final states of two new task, Beam Assembly and Stool Assembly, please refer to Figure 3.
Task Name 5-part Beam Assembly 9-part Stool Assembly
Initial State Refer to caption Refer to caption
Final State Refer to caption Refer to caption
Figure 3: Simulation setup for Beam Assembly and Stool Assembly in IsaacLab.
A.2 Experiment Details
Real Workcell Setup. The real workcell setup includes a UR10e robotic arm and a third-person view Zivid camera (Figure 1). The simulation platform for RL training is built on the IsaacLab engine [51], while motion-level manipulation plans are computed using lazyPRM [61]. Model training is performed on an Ubuntu machine equipped with an Intel i7 CPU and a GeForce RTX 4090 GPU. The Beam Assembly and the Stool Assembly are evaluated in simulation.
Hyperparameters and Computation. For pose estimation, we use $512\times 512$ RGB-D image as input. For high-level policy $\pi_{\theta}$, We adopt the Diffusion Transformer model [53] for predicting action feasibility score. Our RL primitive is trained using 1,000 parallel environments in IsaacLab. The maximum task horizon is 25,000 action steps, equivalent to 200 seconds of robot execution at a 125Hz control frequency.
For our DiT-based high-level policy, the number of reverse diffusion timesteps is a critical parameter that significantly impacts both the sampling time and the quality of the generated samples. We choose 128 diffusion steps, which provide the best balance for most tasks. Additional hyperparameters for the DiT network are listed in Table 3.
A.3 Pose Estimation Results
Accurate pose estimation is essential for enabling high-precision assembly. This is particularly true for the GRASP primitive in our low-level primitives library, where precise object localization directly affects execution success. Figure 4 presents some qualitative pose detection results. The high accuracy of the pose estimator enables us to have a high success rate in terms of both the low-level MP policies and the high-level policy.
Figure 4: Pose Estimation Examples. Our method gives accurate pose estimation of differently shaped parts. The top row shows the input images, and the bottom row shows the pose estimation. Red, green, blue colors indicate the xyzxyz axes in the canonical space, respectively.
A.4 Baseline Description
To evaluate the effectiveness of our proposed approach, we compare it with several adapted hierarchical baselines. However, due to differences in their design, it is challenging to directly apply their methods. In this section, we provide further details about these baselines, including their structure and the adaptations made to align them with our long-horizon assembly tasks.
  • MimicPlay [5]: Since we do not have access to human play data, we retain the hierarchical architecture of MimicPlay but train it using our own expert demonstrations. We modify the high-level policy to predict motion trajectories, and adapt the low-level policy to produce executable actions accordingly. This baseline does not incorporate low-level action primitives.
  • Luo et al. [49]: We adopt the hierarchical structure from their framework. However, for this baseline, we construct the entire low-level primitive library using imitation learning rather than motion planning or reinforcement learning, which leads to primitives that are less robust and adaptable in diverse scenarios.
Table 3: Hyperparameters for high-level policy
Hyper-parameter Value
Hidden Dimension 128
Number of Blocks 4
Number of Heads 4
MLP Ratio 4
Dropout Prob 0.1